OpenStudy (anonymous):

Complete the solution of the equation. Find the value of y when x equals 2. -3x-8y-10

OpenStudy (anonymous):

my guess was y= 0.5

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (whpalmer4):

no need to guess: if you know the value of x, you plug it into the equation and solve for the value of y. Assuming your equation is \[-3x-8y=-10\]Substitute in \(x =2\)\[-3(2) -8y=-10\]\[-6-8y=-10\]Add 6 to both sides\[-8y=-10+6\]\[-8y=-4\]Divide both sides by -8\[y = \frac{1}2\]